# Discovering Effective Policies for Land-Use Planning

### v3

@@ 1. Introduction
One of the main factors contributing to climate change are CO2 emissions due to land-use change (Friedlingstein et al., 2023). Land-use CO2 emissions crucially depend on how much land area is allocated for different uses. Forests in general take up more carbon than e.g. croplands and rangelands, yet such uses are essential for the economy. Land-use patterns must therefore be planned to maximize carbon removals while maintaining economic viability.
@@
Project code and data are available at https://github.com/<anony-mous>, and an interactive demo at https://landuse.evolution.ml.
@@ 3.1. Data
    range: Grazed natural land (i.e., grazing on grassland, savannah, etc.)
@@ 3.2. Decision-making Problem
LinReg (EU) 0.085 0.033 0.171 0.169 0.206
LinReg (SA) 0.570 0.136 0.153 0.062 0.111
LinReg (US) 0.343 0.140 0.146 0.035 0.073
LinReg (Global) 6.872 0.138 0.150 0.036 0.074
RF (EU) 16.493 $0.052$ $0.208^{{\dagger}}$ $0.153^{*{\dagger}}$ $0.220^{{\dagger}}$
RF (SA) 206.451 $0.133^{{\dagger}}$ $\textbf{0.063}^{*{\dagger}}$ $0.076^{{\dagger}}$ $0.127^{{\dagger}}$
RF (US) 101.790 $0.142$ $0.184^{{\dagger}}$ $\textbf{0.019}^{*{\dagger}}$ $0.092^{{\dagger}}$
RF (Global) 422.284 $0.035^{*{\dagger}}$ $0.067^{*{\dagger}}$ $0.020^{*{\dagger}}$ $\textbf{0.040}^{*{\dagger}}$
NeuralNet (EU) 10.650 $\textbf{0.024}^{*{\dagger}}$ $0.266$ $0.282$ $0.333$
NeuralNet (SA) 96.430 $0.252$ $0.099^{*}$ $0.551$ $0.412$
NeuralNet (US) 71.976 $0.127^{*{\dagger}}$ $0.232$ $0.024^{*}$ $0.147$
NeuralNet (Global) 1047.710 $0.045^{*}$ $0.109^{*}$ $0.025^{*}$ $0.048^{*}$
@@
Table 1. Average training time and mean absolute error in tC/ha of the trained models of each type trained on each region and evaluated on that region as well as all other regions. Statistical significance was estimated in each case with the unpaired $t$-test. The RF models are more accurate than NeuralNets (with the exception of EU), which are more accurate than LinReg, and overall the specialized models are more accurate than the Global models. However, RF models do not extrapolate well (as seen in Figure 2), which is necessary for evaluating prescriptor candidates. In contrast, NeuralNets do, and they are also otherwise reasonably accurate. Therefore, the Global NeuralNet model was used to evolve the prescriptors.
@@ 5.1. Prediction
Upon examining the global dataset, a linear relationship was found between ||||ELUC|||| and land-use change with an R-squared value of 0.92. Therefore, a linear regression (LinReg) was fit to the data as the first predictive model. In preliminary experiments, the LinReg model was found to perform the best when current land use, latitude, longitude, cell area, and time were removed from the inputs. Thus, the difference in land use was used as the only input feature in the experiments. Data from the years 1851–2011 (inclusive), totalling 39,209,670 samples globally, was used to train the LinReg models; the 2,420,350 remaining samples from 2012–2021 (inclusive) were used for testing.
@@
Third, a neural network (NeuralNet) was trained for the same task with the full set of features and data from 1851–2011 using the PyTorch library Paszke et al. (2019). The network was fully connected, composed of an input layer, a single hidden layer of size 4096, and an output layer. The output layer took as input the concatenation of the input layer and the hidden layer in order to take advantage of the observed linear relationships more easily (Cheng et al., 2016). It was trained for three epochs with batch size 2048 using the AdamW optimization algorithm Loshchilov and Hutter (2019). After each epoch the learning rate was scaled by a factor of 0.1.
Refer to caption ($a$) LinReg predictions
Refer to caption ($b$) RF predictions
Refer to caption ($c$) NeuralNet predictions
@@
Figure 2. Visualizing the differences in model behavior. Predictions are created for the Global models using synthetic data created by changing 100% of land-use type A (row) to 100% of type B (column) in a 1% sample across the range of latitude, longitude, cell area, and year occurring in the training data. The results are averaged for each conversion from A to B. The models generally agree on the sign of ELUC, suggesting that the results are reliable. The RF model is not able to extrapolate to large values, resulting in low predictions (shown in a much smaller color scale to still make them visible); LinReg and NeuralNet are similar but differ numerically, presumably due to the differences between linear and nonlinear predictions.
@@
LinReg, RF, and NeuralNet models were trained separately on each region and globally. While LinReg is deterministic and needed to be trained only once, RF and NeuralNet are stochastic in their initialization and training and were therefore trained 30 times to evaluate statistical significance. All models were tested on all regions with data from the most recent years, i.e. 2012–2021. As shown in Table 1, the LinReg models were fast to train, but they performed the worst (i.e., they have the largest mean absolute error in almost all cases), suggesting that the linear relationship is not enough to account for the data fully. RF performed significantly better than LinReg and NeuralNet in all regions except Western Europe. However, RF models do not generalize/extrapolate well to large changes in land use that occur in the dataset infrequently (or not at all for some land types; Figure 2), which turns out to be important when evolving prescriptors. Many prescriptor candidates suggest large land-use changes, and the predictor model needs to be able to evaluate them. In contrast, NeuralNets capture nonlinearities, which explains why they outperform LinReg in all regions, as well as RF in Western Europe. Most importantly, they extrapolate well, which suggests that they are a better choice for the surrogate model in prescriptor evolution. Additionally, LinReg and NeuralNet models have many fewer parameters than RF (taking 4KB and 363KB on disk respectively vs. 14GB for RF).
@@
In order to visualize the behavior of the predictor models, they were queried with a synthetic dataset created with every possible conversion from 100% land type A to 100% land type B (the types primf, primn, and urban were excluded from B as was explained in Section 3.2). Because the predictor takes latitude, longitude, cell area, and year as input, 392,096 (1%) latitude/longitude/cell-area/year combinations from the training dataset were sampled randomly and evaluated with each conversion. The results were averaged across the samples to get the overall result for the conversion. Figure 2 shows these results for the Global models. The first finding is that all models generally rank conversions to crops $<$ pasture $<$ forest in terms of reducing ELUC, and generally agree on the sign of ELUC. This consistency suggests that the results are reliable. Second, while LinReg and NeuralNet make similar predictions overall, their numerical values differ in many places, presumably due to the nonlinearities. Third, despite RF’s generally high accuracy in known scenarios, it is not able to generalize/extrapolate to large changes, and instead predicts low-magnitude ELUC values for all imposed land-use transitions.
@@
Thus, since NeuralNet achieves better performance than the RF in most cases, extrapolates well to extremes, and can learn nonlinear relationships in the data, the Global NeuralNet was used as the surrogate model to evolve the prescriptors.
@@ 5.2. Prescriptions
Refer to caption ($a$) Evolution of the Pareto front
Refer to caption ($b$) Performance of all evaluated prescriptors
@@
Refer to caption ($a$) Average performance
Refer to caption ($b$) Performance on underlying samples
Refer to caption ($c$) Performance differences
@@
  • initialization_distribution: orthogonal
@@
As Figure 3 shows, the Evolved Prescriptors and the heuristics all converged in the high-change region of the Pareto front, where 100% forest prescriptions dominate. The heuristics performed slightly better in the low-change region, where the predictors may not be accurate enough to distinguish between small changes. However, in the most useful middle-change region, the Evolved Prescriptors outperformed both heuristics, thus demonstrating that they are likely exploiting the nonlinear relationships well. The performance for each prescription method can be quantified in terms of hypervolume, which is measured as the rectangular area above and to the right of each point (only counting overlapping area once), extending to ELUC = 0 and change = 1. The Even Heuristic is outperformed by the Perfect Heuristic which is outperformed by the Evolved Prescriptors, with hypervolumes of 18.24, 18.87, and 19.13, respectively.
@@ 5.3. Prescriptor Behaviors
Refer to caption ($a$) The location of the cell
Refer to caption ($b$) Suggestions of one Evolved Prescriptor chosen from the Pareto front
@@ 5.4. Interactive Evaluation
As an example, Figure 7 shows the grid cells from which the user has selected one. The system recommends a change in land-use that results in an 18.09 tC/ha reduction in carbon emissions with a 15.77% land-use change. Using the sliders, the user can then explore alternative tradeoffs and modify specific elements of each suggestion. For each variation, the user can then obtain the estimated carbon reduction and land-use change percentage, and in this manner gain an understanding of the possible attainable solutions and confidence in the preferred choice.
@@ 6. Future Work
Another immediate future work direction consists of improving the predictor accuracy through ensembling. Given that models trained on specific regions perform the best on those regions, there is a good chance that an ensemble of them will do even better. Moreover, there may be benefit of including LinReg, RF, and NeuralNet versions of these models in the ensemble as well. There are several ensembling techniques in the Project Resilience codebase, such as averaging, sampling, and confidence-based ensembling, and contributors are encouraged to submit more in the future. The effect of such alternatives will be evaluated in future work, both on prediction accuracy and on driving the search for better prescriptors. The hypervolume of the Pareto front can be used to compare the different approaches quantitatively.
@@ Acknowledgements
We’d like to thank the BLUE, LUH2, and Project Resilience teams, in particular Amir Banifatemi, Prem Krishnamurthy, Gyu Myoung Lee, Gillian Makamara, Michael O’Sullivan, Julia Pongratz, and Jennifer Stave.

### v5

@@ 1. Introduction
One of the main factors contributing to climate change is CO2 emissions due to land-use change (Friedlingstein et al., 2023). Land-use CO2 emissions crucially depend on how much land area is allocated for different uses. Forests in general take up more carbon than e.g. croplands and rangelands, yet such uses are essential for the economy. Land-use patterns must therefore be planned to maximize carbon removals while maintaining economic viability.
@@
Project code and data are available at https://github.com/Project-Resilience/mvp/tree/main/use_cases/eluc, and an interactive demo at https://landuse.evolution.ml.
@@ 3.1. Data
    range: Grazed natural land (i.e., grazing on grassland, savannah, etc.).
@@ 3.2. Decision-making Problem
LinReg (EU) 0.047 0.033 0.172 0.169 0.206
LinReg (SA) 0.457 0.137 0.153 0.061 0.110
LinReg (US) 0.331 0.139 0.146 0.035 0.073
LinReg (Global) 4.644 0.139 0.150 0.035 0.074
RF (EU) 17.697 $0.064$ $0.211^{{\dagger}}$ $0.161^{{\dagger}}$ $0.218^{{\dagger}}$
RF (SA) 209.688 $0.133^{*{\dagger}}$ $\textbf{0.071}^{*{\dagger}}$ $0.074^{{\dagger}}$ $0.126^{{\dagger}}$
RF (US) 111.701 $0.163$ $0.185^{{\dagger}}$ $0.032^{*}$ $0.094^{{\dagger}}$
RF (Global) 417.647 $0.041^{*{\dagger}}$ $0.076^{*{\dagger}}$ $0.028^{*}$ $\textbf{0.045}^{*{\dagger}}$
NeuralNet (EU) 10.711 $\textbf{0.025}^{*{\dagger}}$ $0.277$ $0.286$ $0.334$
NeuralNet (SA) 103.696 $0.248$ $0.100^{*}$ $0.562$ $0.399$
NeuralNet (US) 73.141 $0.136^{{\dagger}}$ $0.225$ $\textbf{0.024}^{*{\dagger}}$ $0.150$
NeuralNet (Global) 1649.193 $0.046^{*}$ $0.110^{*}$ $0.025^{*{\dagger}}$ $0.050^{*}$
@@
Table 1. Average training time and mean absolute error in tC/ha of the trained models of each type trained on each region and evaluated on that region as well as all other regions. Statistical significance was estimated in each case with the unpaired $t$-test. The NeuralNet models are more accurate RF on EU and US, RF more accurate than NeuralNets on SA and globally; both are more accurate than LinReg in all regions (except RF on EU). Overall, the specialized models are more accurate than the Global models. However, RF models do not extrapolate well (as seen in Figure 2), which is necessary for evaluating prescriptor candidates. In contrast, NeuralNets do, and they are also otherwise reasonably accurate. Therefore, the Global NeuralNet model was used to evolve the prescriptors.
@@ 5.1. Prediction
Upon examining the global dataset, a linear relationship was found between ||||ELUC|||| and land-use change with an R-squared value of 0.92. Therefore, a linear regression (LinReg) was fit to the data as the first predictive model. In preliminary experiments, the LinReg model was found to perform the best when current land use, latitude, longitude, cell area, and time were removed from the inputs. Thus, the difference in land use was used as the only input feature in the experiments. Data from the years 1851–2011 (inclusive), totaling 38,967,635 samples globally, was used to train the LinReg models; the 2,420,350 remaining samples from 2012–2021 (inclusive) were used for testing.
@@
Third, a neural network (NeuralNet) was trained for the same task with the full set of features and data from 1851–2011 using the PyTorch library Paszke et al. (2019). The network was fully connected, composed of an input layer, a single hidden layer of size 4096, and an output layer. The output layer took as input the concatenation of the input layer and the hidden layer in order to take advantage of the observed linear relationships more easily (Cheng et al., 2016). It was trained for three epochs with batch size 2048 using the AdamW optimization algorithm Loshchilov and Hutter (2019). After each epoch, the learning rate was scaled by a factor of 0.1.
($a$) LinReg predictions
($b$) RF predictions
($c$) NeuralNet predictions
@@
Figure 2. Visualizing the differences in model behavior. Predictions are created for the Global models using synthetic data created by changing 100% of land-use type A (row) to 100% of type B (column) in a 1% sample across the range of latitude, longitude, cell area, and year occurring in the test data. The results are averaged for each conversion from A to B. The models generally agree on the sign of ELUC, suggesting that the results are reliable. The RF model is not able to extrapolate to large values, resulting in low predictions (shown in a much smaller color scale to still make them visible); LinReg and NeuralNet are similar but differ numerically, presumably due to the differences between linear and nonlinear predictions.
@@
LinReg, RF, and NeuralNet models were trained separately on each region and globally. While LinReg is deterministic and needed to be trained only once, RF and NeuralNet are stochastic in their initialization and training and were therefore trained 30 times to evaluate statistical significance. All models were tested on all regions with data from the most recent years, i.e. 2012–2021. As shown in Table 1, the LinReg models were fast to train, but they performed the worst (i.e., they have the largest mean absolute error in almost all cases), suggesting that the linear relationship is not enough to account for the data fully. RF performed significantly better than LinReg in all regions but EU, significantly worse than NeuralNet in EU and US, and outperformed NeuralNets in SA and globally. However, RF models do not generalize/extrapolate well to large changes in land use that occur in the dataset infrequently (or not at all for some land types; Figure 2), which turns out to be important when evolving prescriptors. Many prescriptor candidates suggest large land-use changes, and the predictor model needs to be able to evaluate them. In contrast, NeuralNets capture nonlinearities, which explains why they outperform LinReg in all regions. Most importantly, they extrapolate well, which suggests that they are a better choice for the surrogate model in prescriptor evolution. Additionally, LinReg and NeuralNet models have many fewer parameters than RF (taking 4KB and 363KB on disk respectively vs. 14GB for RF).
@@
In order to visualize the behavior of the predictor models, they were queried with a synthetic dataset created with every possible conversion from 100% land type A to 100% land type B (the types primf, primn, and urban were excluded from B as was explained in Section 3.2). Because the predictor takes latitude, longitude, cell area, and year as input, 24,204 (1%) latitude/longitude/cell-area/year combinations from the test dataset were sampled randomly and evaluated with each conversion. The results were averaged across the samples to get the overall result for the conversion. Figure 2 shows these results for the Global models. The first finding is that all models generally rank conversions to crops $<$ pasture $<$ forest in terms of reducing ELUC, and generally agree on the sign of ELUC. This consistency suggests that the results are reliable. Second, while LinReg and NeuralNet make similar predictions overall, their numerical values differ in many places, presumably due to the nonlinearities. Third, despite RF’s generally high accuracy in known scenarios, it is not able to generalize/extrapolate to large changes and instead predicts low-magnitude ELUC values for all imposed land-use transitions.
@@
Thus, since NeuralNet achieves similar or better performance than the RF in most cases, extrapolates well to extremes, and can learn nonlinear relationships in the data, the Global NeuralNet was used as the surrogate model to evolve the prescriptors.
@@ 5.2. Prescriptions
($a$) Evolution of the Pareto front
($b$) Performance of all evaluated prescriptors
@@
($a$) Average performance
($b$) Performance on underlying samples
($c$) Performance differences
@@
  • initialization_distribution: orthogonal.
@@
As Figure 3 shows, the Evolved Prescriptors and the heuristics all converged in the high-change region of the Pareto front, where 100% forest prescriptions dominate. The heuristics performed slightly better in the low-change region, where the predictors may not be accurate enough to distinguish between small changes. However, in the most useful middle-change region, the Evolved Prescriptors outperformed both heuristics, thus demonstrating that they are likely exploiting the nonlinear relationships well. The performance for each prescription method can be quantified in terms of hypervolume, which is measured as the rectangular area above and to the right of each point (only counting overlapping area once), extending to ELUC = 0 and change = 1. The Even Heuristic is outperformed by the Perfect Heuristic which is outperformed by the Evolved Prescriptors, with hypervolumes of 19.68, 20.30, and 20.52, respectively.
@@ 5.3. Prescriptor Behaviors
($a$) The location of the cell
($b$) Suggestions of one Evolved Prescriptor chosen from the Pareto front
@@ 5.4. Interactive Evaluation
As an example, Figure 7 shows the grid cells from which the user has selected one. The system recommends a change in land use that results in an 18.09 tC/ha reduction in carbon emissions with a 15.77% land-use change. Using the sliders, the user can then explore alternative tradeoffs and modify specific elements of each suggestion. For each variation, the user can then obtain the estimated carbon reduction and land-use change percentage, and in this manner gain an understanding of the possible attainable solutions and confidence in the preferred choice.
@@ 6. Future Work
Another immediate future work direction consists of improving the predictor accuracy through ensembling. Given that models trained on specific regions perform the best on those regions, there is a good chance that an ensemble of them will do even better. Moreover, there may be a benefit of including LinReg, RF, and NeuralNet versions of these models in the ensemble as well. There are several ensembling techniques in the Project Resilience codebase, such as averaging, sampling, and confidence-based ensembling, and contributors are encouraged to submit more in the future. The effect of such alternatives will be evaluated in future work, both on prediction accuracy and on driving the search for better prescriptors. The hypervolume of the Pareto front can be used to compare the different approaches quantitatively.
@@ Acknowledgements
We would like to thank the BLUE, LUH2, and Project Resilience teams, in particular Amir Banifatemi, Prem Krishnamurthy, Gyu Myoung Lee, Gillian Makamara, Michael O’Sullivan, Julia Pongratz, and Jennifer Stave.